OpenStudy (sanchez9457):

Double Integral:

OpenStudy (sanchez9457):

\[\int\limits_{0}^{1}\int\limits_{0}^{1} xye^xe^y dydx\]

OpenStudy (sanchez9457):

i just can't quite figure this one out! Anyone willing to help?

OpenStudy (kirbykirby):

\[\int\limits_{0}^{1}\int\limits_{0}^{1}xye^xe^ydydx=\int\limits_{0}^{1}ye^ydy \int\limits_{0}^{1}xe^xdx\] Because the "ye^y dy" part doesn't depend on x, it's like a constant so you can tak it out of the integral. Now you have a multiplication of two integrals that can be done by Integration By Parts... your typical xe^x integral :)

OpenStudy (kirbykirby):

The answer will be 1 btw

OpenStudy (tkhunny):

\(\int y\cdot e^{y}\;dy = \int y\;d\left(e^{y}\right) = y\cdot e^{y} - \int e^{y}\;dy\)

OpenStudy (kirbykirby):

Do you need more help

OpenStudy (sanchez9457):

@kirbykirby i really like that answer as you just answered another of my questions!! But i need some clarification if possible:

OpenStudy (sanchez9457):

Okay so basically you just said that the double integral equals the multiple of the integral of each variable?

OpenStudy (kirbykirby):

Yes, In this case you can do it because the term \[ye^ydy\]in the integral doesn't depend on x (recall you can just switch the dydx to dxdy and not worry about the bounds of integration since they are constants and the same for both integrals). So since ye^y dy doesn't depend on x, it acts like a constant (with respect to x), so you can just move it out of the integral

OpenStudy (sanchez9457):

AH!!!!!!!!!!

OpenStudy (sanchez9457):

You @kirbykirby are a genius my good sir!

OpenStudy (kirbykirby):

:) no problem!